OpenStudy (anonymous):

Find the domain and the range of this function. y=125-12x

OpenStudy (rulnick):

domain and range are both "all real numbers"

OpenStudy (anonymous):

what did you do to get that?can you show me?

OpenStudy (rulnick):

unless stated otherwise, domain is all real numbers for which the expression (on the right-hand side) is well defined. this usually just means you have to exclude values that would lead to division by zero. (there are other ways an expression can be undefined.)

OpenStudy (rulnick):

so no matter what x you choose, the RHS is well defined ... so the domain is all reals.

OpenStudy (anonymous):

1. y = 125 - 12x Since the expression is linear, we obtain all real values for y = 125 - 12x. 2. y = 125 - 12x³ Since the degree is odd, the domain and range both consist of all real values

OpenStudy (rulnick):

the range is the set of values that y takes; again, all real numbers. (you can make y be anything you want by choosing the right value of x from the domain.)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Domain and range of linear equation always R